OpenStudy (anonymous):

help? PLEAASE??? What solve or help me understand this: X ---- =-15 -3 I dont get this at all PLEASE help me understand it...im in 8th grade BTW and i need a ton of help please please PLEASE :)

OpenStudy (anonymous):

45 :)

OpenStudy (anonymous):

-3 x -15= 45

OpenStudy (diyadiya):

\[\Large \frac{x}{-3}=-15\] You have to find x So multiply both sides by x \[\Large \frac{x}{-3} \times -3 = -15 \times -3 \]

OpenStudy (anonymous):

@HaperFink22 wow thanks i got an A100% on that one thanks for answering it now i need to understand it :L

OpenStudy (diyadiya):

did you understand?

OpenStudy (anonymous):